OpenStudy (anonymous):

Find the values of k that make 2x^2 + kx + 11 factorable. A. 1 or 11 B. 11 or 13 C. 13 or 23 D. 2 or 11

OpenStudy (mathmate):

K has to be the sum of two integers whose product equals 2*11, as Hero pointed out: ac=2*11=22 k=sum of two integers whose product equals 22.

OpenStudy (anonymous):

Just plug in the values for k, from the given solutions and see which ones work.

OpenStudy (amistre64):

all quadratics are factorable ... is the domain restricted to reals by chance?